# Video-to-Audio Generation with Hidden Alignment

### v1

@@ Abstract
Generating semantically and temporally aligned audio content in accordance with video input has become a focal point for researchers, particularly following the remarkable breakthrough in text-to-video generation. In this work, we aim to offer insights into the video-to-audio generation paradigm, focusing on three crucial aspects: vision encoders, auxiliary embeddings, and data augmentation techniques. Beginning with a foundational model VTA-LDM built on a simple yet surprisingly effective intuition, we explore various vision encoders and auxiliary embeddings through ablation studies. Employing a comprehensive evaluation pipeline that emphasizes generation quality and video-audio synchronization alignment, we demonstrate that our model exhibits state-of-the-art video-to-audio generation capabilities. Furthermore, we provide critical insights into the impact of different data augmentation methods on enhancing the generation framework’s overall capacity. We showcase possibilities to advance the challenge of generating synchronized audio from semantic and temporal perspectives. We hope these insights will serve as a stepping stone toward developing more realistic and accurate audio-visual generation models [^1] The project page: https://sites.google.com/view/vta-ldm.
@@ 1 Introduction
    We propose an effective VTA framework, denoted as VTA-LDM, which achieves SOTA performance in VTA tasks;
@@
    We explore three key aspects within the VTA paradigm: the vision encoder, the auxiliary embedding, and the data augmentation;
@@
    Through a comprehensive evaluation, we offer valuable insights into the model’s design and training process.
@@ 4.1 Dataset
Our primary experiments are conducted on VGGSound Chen et al. (2020), a dataset comprising over 550 hours of videos with acoustic visual-audio event pairs. We train our models on 200k videos, and validate and evaluate them on 3k videos. For certain ablations, we either construct an augmented training set based on the original dataset by randomly cutting and combining segments, or import large unlabelled audio or video corpora for pre-training purposes. More details regarding the dataset preparation, training, and evaluation procedures can be found in section 5.4 and section 5.5.
@@ 4.3 Subjective Evaluation
We include a subjective evaluation of our models, assessing overall quality, audio quality, video-audio semantic alignment, and video-audio temporal alignment using a scale of 1 to 100. Detailed definition of these metrics can be found in the supplementary materials.
@@ 5.2 Initial Results
We first compare our foundational model with other existing VTA baselines to show its capability in audio generation. Specifically, we compare with these open-sourced baselines:
@@
Table 1: Comparison with other baselines.
GT - - - - 0.316 0.492 0.813 0.237
IM2WAV CLIP 6.32 7.46 42.13 2.38 0.182 0.306 0.781 0.192
Diff-Foley CAVP 7.32 9.62 41.09 6.03 0.226 0.441 0.802 0.187
Seeing&Hearing ImageBind 7.32 5.83 32.92 2.62 - - - -
T2AV VA-CLAP 4.05 8.02 33.29 2.12 - - - -
VTA-LDM (Ours) CLIP4Clip 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
Table 1 presents the results of our foundational model compared to existing VTA baselines with different vision encoders (VE), demonstrating its capability in audio generation [^2] For Seeing&Hearing and T2AV, as the codes are not publicly released, we adopt the metrics from their original papers.. We observe that our model outperforms both baselines across all metrics, showcasing its superiority in generating high-quality, diverse, and temporally-aligned audios. Specifically, our model achieves a lower FAD and FD compared to the IM2WAV and Diff-Foley baseline, indicating better distribution similarity between generated and ground truth audios. Regarding CLAP, CAVP, and AV-Align, our model exhibits superior performance compared to the baselines, emphasizing its ability to generate semantically-related and temporally-aligned audios.
5.3 Experiments on Vision Encoders
@@
Table 2: Ablation on vision encoders.
Clip4Clip(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
ImageBind 8.40 6.85 42.18 8.41 0.166 0.186 0.791 0.189
LanguageBind 2.94 9.03 28.65 4.76 0.201 0.338 0.798 0.200
ViViT 15.99 3.88 58.14 7.33 0.288 0.223 0.766 0.189
V-JPEA 14.28 4.27 60.04 11.89 0.142 0.080 0.786 0.186
CAVP 4.27 4.83 56.77 3.88 0.192 0.137 0.787 0.220
5.4 Experiments on Auxiliary Embeddings
@@
Table 3: Ablation on auxiliary embeddings.
Text-only 1.98 10.02 21.25 2.39 0.278 0.486 0.764 0.176
Vision-only(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
Vision+Text 2.09 10.61 21.13 2.47 0.284 0.486 0.802 0.228
Vision+PE 2.36 8.16 29.71 4.88 0.250 0.332 0.801 0.223
Vision+Text+PE 2.99 12.71 25.98 2.76 0.280 0.476 0.794 0.189
Vision+Optical Flow 3.02 9.88 28.32 3.01 0.230 0.473 0.792 0.224
5.5 Experiments on Data Augmentation
@@
Table 4: Ablation on data augmentation.
None(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
Data Clean 3.85 8.19 33.43 0.95 0.314 0.484 0.802 0.281
Concat Augment 2.47 10.82 25.09 3.59 0.236 0.393 0.798 0.248
Pretrain-Video 2.93 9.90 32.10 1.28 0.292 0.267 0.769 0.262
Pretrain-Audio 2.48 12.23 23.61 1.38 0.232 0.469 0.798 0.250
The influence of different data augmentation methods on the performance of the video-to-audio generation model can be observed in table 4. We compare the base model without augmentation to models utilizing various data augmentation techniques. Data Clean leads to the most significant performance improvement. Removing noisy or irrelevant video-audio samples results in improvements mainly in audio quality and video-audio alignment. Concat Augment leads to an improvement in IS, suggesting that this technique can enhance the diversity of the generated audio content, as well as a significant improvement in AV-Align, indicating better video-audio alignment. However, the distance metrics, such as FAD and KL, show a slight decrease with these data augmentation methods. It is worth noting that while data augmentations within the data corpus can improve diversity and alignment, there may be a trade-off in terms of audio content itself. Pretraining with extra data is a natural idea that can help address this issue. The results show that pretraining on extra unlabelled video and audio data can help improve some generation metrics. We believe an appropriate combination of these augmentation methods can enhance the overall performance of the VTA task.
6 Further Experiments & Discussion
6.1 Subjective Evaluation
Table 5: Subjective evaluation results.
GT 91.06$\pm$8.87 91.14$\pm$12.54 89.66$\pm$14.04 91.06$\pm$10.62
IM2WAV 75.01$\pm$16.77 58.35$\pm$28.94 48.33$\pm$20.85 54.45$\pm$21.59
VTA-LDM 72.52$\pm$20.71 64.92$\pm$24.51 43.50$\pm$25.52 49.04$\pm$22.23
Figure 3: Demos of the VTA generation on open-domain videos. Videos are collected from Youtube or generated by OpenAI Sora (Brooks et al., 2024).
6.2 Overall discussion
7 Conclusion
In this paper, we delve into the VTA task, which aims to generate semantically-related and temporally-aligned audios given silent video pieces. Leveraging the diffusion-based backbone of the VTA model, we demonstrate the effectiveness of our VTA-LDM framework and conduct various ablations in the model design and training process. We foresee the evolution of more realistic and natural video-audio generation models based on the insights gained from this study.
Limitations & future work
The models presented in this paper are trained on a limited dataset, VGGSound, which primarily contains audios of single audio event data. Although we employ data augmentations and utilize a portion of the filtered YouTube test set, there is room for improvement. As part of our future work, we plan to build a more extensive, more diverse, and real-world-like high-quality dataset. This will enable us to train and test our models on more complex samples, further enhancing the performance and generalizability of the proposed methods in the VTA domain.
Social impact
The capability to generate realistic audio from silent visual input can significantly enhance the accessibility of high-quality video-audio content, benefiting fields such as artificial intelligence-generated content (AIGC) and virtual reality simulations. However, potential ethical concerns arise, including the possibility of misuse in creating deep fake videos or misleading audio content.

### v2

@@ Abstract
Generating semantically and temporally aligned audio content in accordance with video input has become a focal point for researchers, particularly following the remarkable breakthrough in text-to-video generation. In this work, we aim to offer insights into the video-to-audio generation paradigm, focusing on three crucial aspects: vision encoders, auxiliary embeddings, and data augmentation techniques. Beginning with a foundational model built on a simple yet surprisingly effective intuition, we explore various vision encoders and auxiliary embeddings through ablation studies. Employing a comprehensive evaluation pipeline that emphasizes generation quality and video-audio synchronization alignment, we demonstrate that our model exhibits state-of-the-art video-to-audio generation capabilities. Furthermore, we provide critical insights into the impact of different data augmentation methods on enhancing the generation framework’s overall capacity. We showcase possibilities to advance the challenge of generating synchronized audio from semantic and temporal perspectives. We hope these insights will serve as a stepping stone toward developing more realistic and accurate audio-visual generation models.
@@ 1 Introduction
    We introduce an effective VTA framework, named VTA-LDM, which attains state-of-the-art performance in VTA tasks.
@@
    We investigate three fundamental aspects within the VTA paradigm, providing valuable insights into model design and training processes: 1) a linear combination of semantic video features is sufficient for encoding information required for aligned audio generation; 2) additional auxiliary embeddings assist in filtering out chaotic visual information and focusing on crucial elements; 3) pre-training enhances overall generation capability, and filtering high-quality data improves audio-video alignment.
@@
    Through numerous ablation experiments and evaluations, we supply detailed quantitative results for VTA tasks.
@@ 4.1 Dataset
Our primary experiments are conducted on VGGSound Chen et al. (2020), a dataset comprising over 550 hours of videos with acoustic visual-audio event pairs. We train our models on 200k videos and evaluate them on 3k videos. For certain ablations, we either construct an augmented test set based on the original dataset by randomly cutting and combining segments, or import large unlabelled audio or video corpora for pre-training purposes. More details regarding the dataset preparation, training, and evaluation procedures can be found in section 5.4 and section 5.5.
@@ 4.3 Subjective Evaluation
We include a subjective evaluation of our models, assessing overall quality, audio quality, video-audio semantic alignment, and video-audio temporal alignment using a scale of 1 to 100. In the subjective evaluation, we combine the generated audio with the original video pieces, and show the combined video to the human participants without telling them the concrete test cases. Detailed definition of these metrics can be found in appendix B.
@@ 5.2 Initial Results
We first compare our foundational model with other existing VTA baselines to show its capability in audio generation, including IM2WAV (Sheffer & Adi, 2023), Diff-Foley (Luo et al., 2024), FoleyCrafter (Zhang et al., 2024), Seeing&Hearing (Xing et al., 2024) and T2AV (Mo et al., 2024). These compared baselines primarily utilize the diffusion framework. Although there are additional VTA baselines available, we consider these to be representative and indicative of SOTA performance. We also list the vision encoders these baselines use for reference. For certain benchmarks, we omit their alignment metrics, as they are tailored to produce shorter audio segments, which does not align with competitive standards. See appendix A for more details about these baselines.
Table 1: Comparison with other baselines.
GT - - - - 0.316 0.492 0.813 0.237
IM2WAV CLIP 6.32 7.46 42.13 2.38 0.182 0.306 0.781 0.192
Diff-Foley CAVP 7.32 9.62 41.09 6.03 0.226 0.441 0.802 0.187
FoleyCrafter CLIP 4.44 9.44 27.00 4.57 0.307 0.179 0.762 0.239
Seeing&Hearing ImageBind 7.32 5.83 32.92 2.62 - - - -
T2AV VA-CLAP 4.05 8.02 33.29 2.12 - - - -
VTA-LDM (Ours) CLIP4Clip 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.243
Table 1 presents the results of our foundational model compared to existing VTA baselines with different vision encoders (VE), demonstrating its capability in audio generation [^1] For Seeing&Hearing and T2AV, as the codes are not publicly released, we adopt the metrics from their original papers.. We observe that our model is competitive with all of the baselines across the metrics, showcasing its superiority in generating high-quality, diverse, and temporally-aligned audios. Specifically, our model achieves a lower FAD and FD compared to the IM2WAV and Diff-Foley baseline, indicating better distribution similarity between generated and ground truth audios. Regarding CLAP, CAVP, and AV-Align, our model exhibits superior performance compared to the baselines, emphasizing its ability to generate semantically-related and temporally-aligned audios.
5.3 Experiments on Vision Encoders
Table 2: Ablation on vision encoders.
Clip4Clip(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
ImageBind 8.40 6.85 42.18 8.41 0.166 0.186 0.791 0.189
LanguageBind 2.94 9.03 28.65 4.76 0.201 0.338 0.798 0.200
ViViT 15.99 3.88 58.14 7.33 0.288 0.223 0.766 0.189
V-JPEA 14.28 4.27 60.04 11.89 0.142 0.080 0.786 0.186
CAVP 4.27 4.83 56.77 3.88 0.192 0.137 0.787 0.220
We begin by investigating how different vision encoders may influence the VTA results. Our study includes popular vision encoders such as Clip4Clip (Luo et al., 2021), Imagebind (Girdhar et al., 2023), LanguageBind (Zhu et al., 2023), V-JEPA (Bardes et al., 2024), ViViT (Arnab et al., 2021), and CAVP (Luo et al., 2024). These vision encoders can primarily be categorized into two groups: 1) image encoders trained with multimodal alignment to enhance semantic understanding (e.g., Clip4Clip and Imagebind), and 2) video feature encoders designed to extract semantic and temporal information from videos. Details can be found in appendix C.
Figure 2: The saliency map of our model’s interest upon the visual input. We illustrate that VTA-LDM has the ability to learn and concentrate on potential objects capable of producing sound. Furthermore, the model is designed to focus on various sections of the frame across different time intervals, although the attention is calculated based on the final audio latent representation only once.
To further demonstrate how the model’s generation part engagements with the input video, we compute the saliency map of the input visual feature based on the generated audio content, aiming to empirically show how VTA-LDM processes the input video frames. As illustrated in fig. 2, our model focuses on different potential video regions that could potentially be the source of the audio. The model can engage in an attention shuffle with the frames as they progress, indicating its ability to capture temporal information between the frames, despite using the pure semantic vision encoder CLIP. For a safe conclusion, we note that a simple temporal combination of semantic vision features of the video frames may be sufficient to capture the features from the video in VTA tasks, eliminating the need for re-training a video feature encoder.
5.4 Experiments on Auxiliary Embeddings
Auxiliary embeddings serve as additional information beyond visual features, potentially enhancing video understanding and anchoring audio events to improve the generation process. We explore the integration of extra information from various modalities, including textual descriptions that encapsulate the overall video content, position embeddings that indicate the sequence of frames, and optical flow that captures pixel-level changes between frames within the video. Details can be found in appendix D.
Table 3: Ablation on auxiliary embeddings.
Text-only 1.98 10.02 21.25 2.39 0.278 0.486 0.764 0.176
Vision-only(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
Vision+Text 2.09 10.61 21.13 2.47 0.284 0.486 0.802 0.228
Vision+PE 2.36 8.16 29.71 4.88 0.250 0.332 0.801 0.223
Vision+Text+PE 2.99 12.71 25.98 2.76 0.280 0.476 0.794 0.189
Vision+Optical Flow 3.02 9.88 28.32 3.01 0.230 0.473 0.792 0.224
Figure 3: A Comparison Between Models Without Additional Text Embedding. The left saliency maps encode the text embedding, while the right ones do not. We demonstrate that extra text embeddings can aid the model in gaining a deeper understanding of the visual content.
We observe that additional embeddings, particularly extra text embeddings, can not only assist the model in comprehending the input video but also concentrate on the actual object producing the audio, thereby improving audio-visual alignment, as illustrated in fig. 3. Visual inputs are more complex and chaotic, making it challenging to discern the true source of the audio. For instance, (on the top left) a woman speaks without moving her mouth, but a non-fine-grained vision encoder cannot recognize this. The extra textual input can help identify the real focus, which is the ”guitar.”
5.5 Experiments on Data Augmentation
Table 4: Ablation on data augmentation.
None(VTA-LDM) 2.05 10.10 25.50 3.81 0.245 0.452 0.800 0.225
Data Alignment Filter 3.85 8.19 33.43 0.95 0.314 0.484 0.802 0.281
Concat Augment 2.47 10.82 25.09 3.59 0.236 0.393 0.798 0.248
Pretrain-Video 2.93 9.90 32.10 1.28 0.292 0.267 0.769 0.262
Pretrain-Audio 2.48 12.23 23.61 1.38 0.232 0.469 0.798 0.250
The data used in training is also crucial. We explore several approaches from different perspectives. Initially, we filter the training set using video-audio alignment labels to obtain high-quality video-audio pairs. We use a CLAP (Wu* et al., 2023) model to help select audio-video pairs with similar semantics based on extra textual labels (with $score>0.3$). We also use the AV-Align score to filter unmatched video-audio pairs (with $score>0.2$). Furthermore, we utilize data augmentation techniques to enhance the diversity and complexity of the training set. We randomly combine video and audio segments to construct a diverse and multi-content set of training data points. Additionally, we leverage extra training sets for pretraining and perform fine-tuning on the foundational model. These approaches are separately studied to illustrate their influence on the final generation results. More details can be found in appendix E.
The influence of different data augmentation methods on the performance of the VTA generation model can be observed in table 4. We compare the base model without augmentation to models utilizing various data augmentation techniques. Data Clean leads to the most significant performance improvement. Removing noisy or irrelevant video-audio samples results in improvements mainly in audio quality and video-audio alignment. The drop in generation quality can be attributed to the decrease in the number of training instances. Concat Augment leads to an improvement in IS, suggesting that this technique can enhance the diversity of the generated audio content, as well as a significant improvement in AV-Align, indicating better video-audio alignment. However, the distance metrics, such as FAD and KL, show a slight decrease with these data augmentation methods. It is worth noting that while data augmentations within the data corpus can improve diversity and alignment, there may be a trade-off in terms of audio content itself. Pretraining with extra data is a natural idea that can help address this issue. The results show that pretraining on extra unlabelled video and audio data can help improve some generation metrics. We believe an appropriate combination of these augmentation methods can enhance the overall performance of the VTA task.
6 Further Experiments & Discussion
6.1 Subjective Evaluation
Table 5: Subjective evaluation results.
GT 91.06$\pm$8.87 91.14$\pm$12.54 89.66$\pm$14.04 91.06$\pm$10.62
IM2WAV 75.01$\pm$16.77 58.35$\pm$28.94 48.33$\pm$20.85 54.45$\pm$21.59
VTA-LDM 72.52$\pm$20.71 64.92$\pm$24.51 43.50$\pm$25.52 49.04$\pm$22.23
Figure 5: Demos of the VTA generation on open-domain videos. Videos are collected from YouTube or generated by OpenAI Sora (Brooks et al., 2024). Although some test data points may exhibit styles distinct from those in the training dataset, our model concentrates on the semantic comprehension of video content and possesses a certain degree of out-of-domain generalization capability.
6.2 Overall discussion
7 Conclusion
In this paper, we delve into the VTA task, which aims to generate semantically-related and temporally-aligned audios given silent video pieces. Leveraging the diffusion-based backbone of the VTA model, we demonstrate the effectiveness of our VTA-LDM framework and conduct various ablations in the model design and training process. We also dive into several key aspects during the VTA generation, aiming at giving the community more insights about the model design and training process. Specifically, we concentrate on the vision encoder, auxiliary embeddings, and data augmentation techniques employed in the training process, and propose several significant suggestions that have been quantitatively validated through our experiments. A combination of these approaches can lead to a more powerful VTA backbone. We foresee the evolution of more realistic and natural video-audio generation models based on the insights gained from this study.
Limitations & future work The models presented in this paper are trained on a limited dataset, VGGSound, which primarily contains audios of single audio event data. Although we employ data augmentations and utilize a portion of the filtered YouTube test set, there is room for improvement. As part of our future work, we plan to build a more extensive, more diverse, and real-world-like high-quality dataset.
Social impact The capability to generate realistic audio from silent visual input can significantly enhance the accessibility of high-quality video-audio content, benefiting fields such as AI-generated content (AIGC) and virtual reality simulations. However, potential ethical concerns arise, including the possibility of misuse in creating deep fake videos or misleading audio content.
Appendix A Baselines
We compare with these open-sourced VTA baselines:
@@ Appendix A Baselines
    FoleyCrafter (Zhang et al., 2024) imports semantic and temporal blocks for precise audio-video synchronization, and supports the use of text descriptions to facilitate controllable and diverse TTA generation.
  4. 4.
@@
Appendix B Subjection Evaluation
The subjective evaluation aims to gauge the performance of our models from a human perspective, providing insights into the perceived quality and alignment of the generated audio and original video content. Participants undergo a training session where they are introduced to the evaluation objectives, the rating scale, and example demonstrations that illustrate different quality levels. Additionally, they practice rating sample videos to familiarize themselves with the criteria and receive feedback to ensure consistency.
Overall Quality
This metric evaluates the general appeal and coherence of the combined video, considering both audio and visual components. It contains how well the audio and video elements fit together without any jarring or inconsistent moments.
Audio Quality
Focusing solely on the generated audio, this metric assesses factors such as clarity, fidelity, and naturalness, like how clear and understandable the audio is, free from distortion or muddiness, and how closely the generated audio resembles human speech or natural sounds in terms of intonation, rhythm, and expression.
Video-Audio Semantic Alignment
This metric measures how well the audio semantically matches the visual content, ensuring that the sounds correspond appropriately to the actions and scenes depicted in the video, like how relevant the audio is to the visual context, with sounds matching the actions and scenes on screen.
Video-Audio Temporal Alignment
The evaluation centers on the synchronization between the audio and video streams, determining how accurately the timing of audio events aligns with visual events, like the precision with which sounds occur in tandem with corresponding visual actions (e.g., a door slamming in sync with the sound of the slam).
Appendix C Experiments on Vision Encoders
@@
Appendix D Experiments on Auxiliary Embeddings
@@
Appendix E Experiments on Data Augmentation
@@
Appendix F More Demos
F.1 Sailency Map
Figure 6: More demos of the sailency map. We show that the model focuses on different components in different frames of the given video after training, although the vision encoder parameters are completely freezed during the training time.
F.2 Audio Generation
Figure 7: More demos of the VTA generation on open-domain videos. Refer to the supp materials for more video demos and the inference code.